# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

### v2

@@ Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions (and, optionally, a list of simple facts for counterexample (CE) construction), written in a minimalist Mathematical Programming Language (MPL), and systematically explores their deductive neighborhood. Definitions are compiled into a distributed grid of simple Logic Blocks (LBs) that exchange messages; whenever the premises of an inference rule unify, a new fact is emitted with full provenance to its sources, yielding replayable, auditable proof graphs. A prototype software implementation instantiates the workflow on first-order Peano arithmetic. Starting only from the Peano axioms, GL enumerates conjectures, applies normalization, type, and CE filter, and automatically reconstructs machine-checkable proofs of foundational arithmetic laws, including associativity and commutativity of addition, associativity and commutativity of multiplication, and distributivity. On commodity hardware, the prover phase requires approximately 7 seconds; a complete run finishes in about 5 minutes. Generated proofs export to navigable HTML so that every inference step can be inspected independently. We outline a hardware-software co-design path toward massively parallel realizations and describe prospective integration with probabilistic models (e.g., large language models) for auto-formalization and conjecture seeding. The Python, C++, and MPL code to reproduce the Peano experiments, along with the full proof graphs in HTML as well as machine-readable text format, are available in the project’s GitHub repository at github.com/Generative-Logic/GL commit 56c9233 and are permanently archived at doi:10.5281/zenodo.17206386.
@@ 1 Introduction
This paper introduces Generative Logic (GL), a novel computer architecture conceived to bridge this gap. GL is designed to automate deep, definition-based mathematical reasoning by shifting the paradigm from human-assisted, single-theorem proving to the fully automatic generation and verification of entire families of theorems. At its core, the system takes a set of formal axioms and algorithmically explores the complete deductive space they define. This approach enables GL to systematically generate conjecture, produce machine-verifiable proofs, and iterate through complex logical chains at computational speeds.
@@
This paper presents the foundational principles, architecture, and early results of GL. We detail the system’s massively parallel, hardware-aware proof engine, its custom Mathematical Programming Language (MPL) for formalizing definitions, and its unique execution model. A case study in Peano arithmetic demonstrates the system’s ability to autonomously derive and prove foundational theorems, validating the architecture and its approach to automated discovery.
@@ 2 The GL Workflow
    Compilation and Disintegration. The generated conjectures are then compiled and distributed across the grid of Logic Blocks (LB). Each Logical Entity (LE) within a conjecture is assigned to a specific block, forming a distributed processing chain. Within each block, the relevant definitions are disintegrated into their constituent logical parts, which serve as the initial fuel for the proof search.
@@ 3.6.3 Reasoning Contexts
This core execution procedure is versatile and can be deployed in different contexts to achieve various reasoning goals. The context determines how the output of the procedure is interpreted. For a proof by contradiction, the system uses the following approach: Assume a conjecture $A\implies(B\implies(C\implies D))$. This is mapped to a processing chain $A\to B\to C$. To prove it by contradiction, the system attaches the negated conclusion, $\neg D$, as a fourth block, creating the chain $A\to B\to C\to\neg D$. This final block is "primed" to detect a contradiction. If, during its execution, the $\neg D$ block generates or receives an expression that is the logical opposite of an expression it already holds (e.g., it derives $P(x)$ when it already possesses $\neg P(x)$), it signals a contradiction LB C. Upon receiving this signal, the LB C the original theorem, $A\implies(B\implies(C\implies D))$, as a new, globally valid theorem and broadcasts it to all other LBs. The same core execution procedure can be similarly adapted for other strategies, such as proof by induction or concrete symbolic calculations.
@@ 3.7.2 The Interactive Proof Graph
Figure 5: A screenshot of a proof chapter (e.g., chapter17.html), showing several inference steps.
@@
Table 1: End-to-end RT by phase on commodity hardware (32 logical cores).
Phase Time (s) Share (%)
Conjecturer 250 87.1
CE filter 30 10.5
Prover 7 2.4
Total 287 100
4 Case Study: Peano Arithmetic
We evaluate GL on first-order Peano arithmetic on commodity hardware (Dell G16 7630, 32 logical cores, 32 GB RAM). The end-to-end wall clock is about five minutes. The three phases contribute as follows: the conjecturer takes roughly 250 seconds, the CE filter about 30 seconds, and the prover about 7 seconds. The prover now runs at interactive speed; overall RT is dominated by conjecture generation and CE screening.
Rationale and headroom.
We deliberately defer aggressive micro-optimizations of the Peano pipeline. Many techniques can overfit to Peano’s shape and risk being useless or harmful on larger theories. Instead, we will pursue optimizations that transfer to Elementary Number Theory (ENT) before investing in phase-specific tuning. This preserves external validity while leaving substantial headroom to reduce the end-to-end RT.
What the numbers say.
  • Bottleneck is the conjecturer (about 87%). Transfer from Python to C++ is the easiest way to drive the current RT down due to stronger multithreading in the latter language. In practice, accelerating the conjecturer has never been a problem so far, taking the smallest amount of development time compared with the prover.
  • CE filter is modest but meaningful (about 10%). ENT-oriented speedups include batched evaluation, RT-friendly reformulation of conjectures, and reducing the amount of simple facts provided to the filter.
  • Prover is no longer the limit (about 2%). We expect a significant blow-up of prover RT for ENT. Ultimately, it depends on whether most wrong conjectures can be filtered by CEs. If, after all layers of aggressive filtering, there remain, for example, 400 mostly correct conjectures, RT should stay under control even with trichotomy.
Takeaway. With a prover RT of approximately 7 seconds and peak memory below a gigabyte, GL’s current bottleneck is conjecture generation. We will drive the remaining RT down during the transition to ENT, prioritizing optimizations (prover runs with tiered parametrization, batched CE tiers) that generalize to larger theories.
The present RT distribution should be read as a snapshot of the current implementation rather than a stable characteristic of GL; as the system and target theories evolve, the relative weights may shift in ways that are hard to predict, though the prover will likely remain the principal bottleneck. Given the many yet-unexplored but plausible optimization avenues, we are moderately optimistic and eager to engage.
@@ 5 Related Work and Contributions
Systems such as Lean, Rocq/Coq, and Isabelle/HOL provide expressive foundations for formalizing mathematics and verifying proofs in a small, trusted kernel [6]. Although their traditional workflow is interactive—users decompose goals, invoke tactics, and supply key intermediate lemmas—modern ecosystems expose substantial automation. Isabelle’s Sledgehammer tool, for example, exports the current goal together with a heuristic relevance slice of the local theory context to a portfolio of external automated theorem provers (ATPs) and SMT solvers (e.g., E, Vampire, Z3, cvc5); when one succeeds, Isabelle attempts proof reconstruction automatically [3]. Lean and Coq likewise support tactic automation and scripted, batch replay of large libraries. In short, contemporary ITPs can discharge many individual goals automatically once the goals and supporting libraries are in place. Coq’s ecosystem has recently incorporated “hammer” plugins (e.g. CoqHammer) that automatically attempt to prove goals by translating them to first-order logic and calling external provers [4] , often discharging nontrivial goals without manual guidance. Automated first-order provers like E have achieved remarkable success in solving individual conjectures, consistently ranking among the top systems in the annual CASC competitions[10]. Another premier first-order prover is Vampire, which has dominated many competition divisions – since 1999 it has won over 50 category titles in the “world cup” of theorem provers [7] – thanks to its highly optimized saturation engine. SMT solvers (e.g. Z3) excel at deciding logical formulas in rich theories automatically [5] , and they have become integral to many verification workflows (often being invoked behind the scenes by proof assistant tactics to discharge specific subgoals). The newest SMT solvers, such as cvc5, extend the capabilities of their predecessors (e.g. CVC4 and Z3) with support for a diverse range of theories and advanced features like higher-order reasoning and syntax-guided synthesis [2].
@@ 6 Future Work, Roadmap and Potential Impact
The successful application of GL to Peano arithmetic serves as a crucial proof of concept, validating the core architecture and its approach to automated reasoning. However, this is merely the first step. The ultimate vision for GL is to tackle progressively more complex mathematical domains. This section outlines the next major milestone for the project and the concrete steps required to achieve it.
6.1 Next Milestone: ENT
The next grand challenge for GL is the automatic formulation of ENT. The goal is to provide the system with the necessary foundational definitions and have it autonomously generate a complete, verifiable proof graph for all major theorems up to and including the Fundamental Theorem of Arithmetic (FTA). Achieving this would represent a significant leap from the current baseline, demonstrating that the GL architecture is not limited to a single, simple theory but is a general-purpose engine for mathematical exploration.
6.2 The Path to Scalability and Performance
The roadmap to achieving the required performance involves a three-stage optimization and scaling plan:
  1. 1.
    Codebase Migration: To further enhance performance, the entire prototype will be migrated from Python to C++. This will provide significant speed improvements due to C++’s lower-level memory management and more efficient execution.
  2. 2.
    Massive Parallelization: The final and most critical step is to fully realize the parallel nature of the GL architecture. It will be adapted to run on massively parallel cloud infrastructure, distributing the workload across thousands of spot instances on platforms like Amazon Web Services (AWS) or Google Cloud Platform (GCP).
6.3 Architectural Enhancements for Advanced Reasoning
Beyond performance, tackling ENT requires several key enhancements to the core capabilities of the GL engine itself. The following architectural upgrades are planned:
Definition Integration
Currently, GL is proficient at disintegrating existing LEs into their constituent parts for use in proofs. The system must be extended with the inverse capability: the ability to integrate or assemble new, complex LEs from available building blocks, such as creating a LE for inequality from the LE of sum and concept of existence.
Introduction of Negation
The theorem weaving process will be enhanced to include negation. This will allow the system to automatically formulate a much richer set of conjectures, including those of the form $\neg A\implies B$ or $A\implies\neg B$.
Implementation of Trichotomy
The proof process will be augmented with the law of trichotomy, allowing the system to reason by cases (e.g., for any two numbers $a$ and $b$, either $a<b$, $a=b$, or $a>b$). This is an essential tool for many proofs in number theory.
By systematically executing this roadmap of performance optimizations and architectural enhancements, we anticipate that GL will be equipped to cover the domain of ENT within a reasonable RT, marking a major step towards the goal of fully automated, large-scale mathematical discovery.

### v3

@@ Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions, written in a minimalist Mathematical Programming Language (MPL), and systematically explores a configurable region of their deductive neighborhood. A defining feature of the architecture is its unified hash-based inference engine, which executes both algebraic manipulations and deterministic logical transformations. Definitions are compiled into a distributed grid of simple Logic Blocks (LBs) that exchange messages; whenever the premises of an inference rule unify, a new fact is emitted with full provenance to its sources, yielding replayable, auditable proof graphs. Experimental validation is performed on Elementary Number Theory (ENT) utilizing a batched execution strategy. Starting from foundational axioms and definitions, the system first develops first-order Peano arithmetic, which is subsequently applied to autonomously derive and prove Gauss’s summation formula as a main result. To manage combinatorial explosion, GL algorithmically enumerates conjectures and applies normalization, type constraints, and counterexample (CE) filtering. On commodity hardware, an end-to-end run completes in under 7 minutes. Generated proofs export as navigable HTML so that every inference step can be inspected independently. We outline a hardware-software co-design path toward massively parallel realizations and describe future integration with large language models (LLMs) for auto-formalization and conjecture seeding. The Python, C++, and MPL code to reproduce these experiments, along with the full proof graphs in HTML as well as machine-readable text format, are available in the project’s GitHub repository at github.com/Generative-Logic/GL commit 1771330 and are permanently archived at doi:10.5281/zenodo.17206386.
@@ 1 Introduction
This paper introduces Generative Logic (GL), a novel computer architecture conceived to bridge this gap. GL is designed to automate deep, definition-based mathematical reasoning by shifting the paradigm from human-assisted, single-theorem proving to the fully automatic generation and verification of entire families of theorems. At its core, the system takes a set of formal axioms and algorithmically explores the constrained deductive space they define. This approach enables GL to systematically generate conjectures, produce machine-verifiable proofs, and iterate through complex logical chains at computational speeds.
@@
This paper presents the foundational principles, architecture, and early results of GL. We detail the system’s massively parallel, hardware-aware proof engine, its custom Mathematical Programming Language (MPL) for formalizing definitions, and its unique execution model. A case study in Elementary Number Theory (ENT) demonstrates the system’s ability to autonomously derive and prove foundational theorems, validating the architecture and its approach to automated discovery.
In this light, GL can be understood through an analogy with biological development. A living organism does not store an explicit blueprint of every cell, tissue, and organ it will produce. Instead, a compact set of genetic instructions — DNA — encodes a small collection of foundational rules, and the machinery of the cell does the rest: molecules interact, differentiate, and self-organize into structures of extraordinary complexity. Critically, this unfolding is not unbounded. Gene regulation, energy budgets, and environmental signals determine which developmental paths are expressed in a given context; the genome encodes far more potential than any single organism realizes. GL operates on the same principle, but in the domain of logic. A minimal set of axiomatic definitions, written in MPL, plays the role of the genome: a compact, high-density encoding of foundational truth. The GL architecture then acts as the developmental machinery — a deterministic engine that explores the deductive consequences those definitions imply, with per-batch configuration files playing a role analogous to gene regulation: they do not alter the logical content of the definitions, but they determine which region of the deductive space is explored in a given run. Just as no biologist needs to manually specify every protein interaction for an organism to grow, no mathematician needs to manually state every theorem for GL to discover it. The axioms are planted; the proof graph grows. As computational resources scale, constraints can be progressively relaxed, allowing broader exploration from the same definitional base. This shift — from explicit enumeration to generative unfolding — is the central architectural commitment of GL, and it is what distinguishes it from both interactive proof assistants, which require human-directed goal-setting, and from probabilistic language models, which approximate reasoning without grounding it in axiomatic truth.
@@ 2 The GL Workflow
The GL pipeline is executed in discrete, ordered batches, where each batch corresponds to a self-contained mathematical domain or theory layer. This batched design serves a primary practical purpose: by demixing domains, it prevents the conjecturer and prover from operating over a conflated search space, which would combinatorially explode as the number of definitions grows. Instead, each batch operates on a focused, well-scoped slice of the theory, with its results becoming available as proven ground for subsequent batches. Each batch is governed by a dedicated configuration file, which exercises two key controls: it defines the search space of the conjecturer — specifying which definitions, operators, and structural templates are eligible for conjecture weaving — and it sets the operational depth of the prover, determining how far the inference engine will chase logical consequences before halting. This separation of concerns between batches, combined with per-batch configurability, is what allows GL to scale incrementally toward richer theories without sacrificing the determinism and auditability that define the architecture.
@@
    Compilation and Disintegration. The generated conjectures are then compiled and distributed across the grid of Logic Blocks (LBs). Each Logical Entity (LE) within a conjecture is assigned to a specific block, forming a distributed processing chain. Within each block, the relevant definitions are disintegrated into their constituent logical parts, which serve as the initial fuel for the proof search.
@@
The natural habitat of GL is a cloud of specialized ASIC chips — a massively parallel substrate where multitudes of Logic Blocks operate simultaneously and the architecture’s full potential is realized without constraint. For the time being, however, GL must survive in the bonsai / resource-constrained conditions of a single machine or a cloud fragment. Configuration files are precisely the instrument that makes this possible. By tightening the conjecturer’s search space and capping the prover’s operational depth, they allow available computational power to be traded against domain-specific knowledge: the less compute one has, the more the configuration must compensate by encoding structural guidance, type constraints, and targeted conjecture templates. Conversely, as hardware scales up, constraints can be relaxed and the system allowed to explore more freely. This mechanism is what tames a beast built for unlimited compute and lets it run, usefully, on a local machine or cloud fragment. The smallest but still meaningful deployment of GL is one where the user supplies not only definitions but all necessary lemmas alongside them, leaving GL the focused task of assembling these ingredients into a verified, navigable proof graph — sacrificing breadth of discovery in exchange for guaranteed tractability.
In the current implementation, simple facts such as addition and multiplication tables are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact — for example, that 3 × 4 = 12 — is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level arithmetic facts, effectively allowing the system to bootstrap its own fact base. However, GL currently lacks case differentiation — the ability to reason by cases such as trichotomy — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts.
@@ 3.4 CE Filter
The CE filter is as a rejection mechanism relative to the supplied ground-fact tables: if a contradiction is found, the conjecture is discarded. However, it is incomplete as a decision procedure: surviving the CE stage does not imply truth or provability, only compatibility with the tested finite facts.
@@
Upcoming extension - Incubator.
In the current implementation, simple facts used by the CE filter are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery, regardless of the domain under exploration. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level facts, effectively allowing the system to bootstrap its own fact base. This tool becomes increasingly necessary as the complexity of the target domain grows: what counts as a "simple" fact in elementary arithmetic is a straightforward table entry, but in richer domains, the corresponding ground-level facts can themselves be non-trivial statements requiring structured derivation. Without automation, the manual curation of such facts quickly becomes a bottleneck. However, GL currently lacks case differentiation — the ability to reason by cases — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts, across any domain expressible in MPL.
@@ 3.6.2 Core Execution Procedure
3.6.3 Deterministic Logical Transformer
The same hash-based inference engine that drives proof search through the application of previously proven theorems — that is, algebraic manipulation — is also responsible for a fundamentally different but equally central task: the deterministic transformation of existing logical entities into new ones, a capability we term the Deterministic Logical Transformer. (The term ’transformer’ is used here in its classical sense — a mechanism that converts one structure into another — and bears no relation to the neural Transformer architecture of Vaswani et al.)
The transformation process consists of two complementary operations, both introduced earlier in the pipeline: disintegration and integration. Disintegration decomposes a compound logical entity into its constituent parts; integration reassembles available parts into new, valid compound entities. Crucially, both operations are governed by the same underlying artifact: the GL binary, the principal output of the compilation phase. The binary encodes exact instructions for how any compound logical structure may be disassembled and how its components may be recombined. The compiled representation is therefore not directional — the same map can be traversed in both ways, making disintegration and integration two faces of a single deterministic process rather than separate mechanisms.
Figure 4 illustrates this structure for the fold definition. The root node is of type existence, which expands into two children: interval (type and) and existence2 (type existence). Each of these expands further into its constituent elements. Nodes of type implication are terminal — they are not expanded but stored as inference rules in the local hash table of the corresponding Logic Block. Predicate leaves represent ground logical statements that serve directly as facts or hash keys during execution. The same graph that guides disintegration — breaking fold down to its atomic parts — equally guides integration, assembling a valid fold instance from available components by traversing the map in reverse.
The roles of the three node categories differ fundamentally between the two directions of traversal. During disintegration, and and existence nodes act as predicate suppliers: their constituent expressions are extracted and injected into the hash engine as known facts, fueling the inference process. Implication nodes, by contrast, are loaded directly as key-value pairs into the hash table, where they serve as reusable inference rules available to the engine at any point during execution. During integration, the roles shift. Implication nodes now become sub-theorems that must be actively proved — the hash engine is tasked with establishing them before the compound entity can be assembled. Meanwhile, and and existence nodes undergo a structural inversion: they are themselves transformed into implications, reframing the question of constructing a compound entity as a conditional one. This allows the hash engine to reason about whether the necessary components are present and, if so, to conclude that the compound entity is validly constructible. Disintegration and integration thus engage the same underlying machinery in opposite modes — one dissolving structure into fuel, the other building structure from proof.
Figure 4: Compiled structure of the fold definition. Existence and and-nodes expand further; implication nodes are terminal stop points; predicate leaves are ground facts. The same map is traversed in both directions by the Deterministic Logical Transformer.
@@ 3.6.4 Reasoning Contexts
This core execution procedure is versatile and can be deployed in different contexts to achieve various reasoning goals. The context determines how the output of the procedure is interpreted. For a proof by contradiction, the system uses the following approach: Assume a conjecture $A\implies(B\implies(C\implies D))$. This is mapped to a processing chain $A\to B\to C$. To prove it by contradiction, the system attaches the negated conclusion, $\neg D$, as a fourth block, creating the chain $A\to B\to C\to\neg D$. This final block is "primed" to detect a contradiction. If, during its execution, the $\neg D$ block generates or receives an expression that is the logical opposite of an expression it already holds (e.g., it derives $P(x)$ when it already possesses $\neg P(x)$), it signals a contradiction LB C. Upon receiving this signal, the LB C emits the original theorem, $A\implies(B\implies(C\implies D))$, as a new, globally valid theorem and broadcasts it to all other LBs. The same core execution procedure can be similarly adapted for other strategies, such as proof by induction or concrete symbolic calculations.
@@ 3.7.2 The Interactive Proof Graph
Figure 6: A screenshot of a proof chapter (chapter81.html), showing several inference steps.
4 Case Study: Elementary Number Theory
To assess the GL system, we conducted a two-stage evaluation encompassing both Peano arithmetic and Gauss’s summation formula. The experiments were performed on standard consumer hardware (a Dell G16 7630 equipped with 32 logical cores and 32 GB of RAM). The complete execution finished in roughly 407 seconds, or just under seven minutes. A detailed breakdown of the time spent in each phase across the two batches is provided in Table 1. Notably, rather than the standard fractional form, GL autonomously derives the division-free variant of Gauss’s formula, expressing the product of $n$ and its successor as twice the sum:
$n(n+1)=2\sum_{i=1}^{n}i$ (1)
Batch 1: Peano Arithmetic
The conjecturer produces 406 candidates in 21 seconds. The CE filter processes all 406 conjectures in 14.6 seconds, passing 55 survivors to the prover. The prover runs 17 iterations (2 warm-up, 15 main) and finishes in 24.6 seconds, establishing 32 theorems. These cover the foundational arithmetic laws of Peano arithmetic, including commutativity and associativity of addition and multiplication, and distributivity. The 64 theorem entries written to file represent mirrored pairs — each theorem together with its symmetric counterpart. The total runtime of the native executable for this batch is 39.4 seconds.
Batch 2: Gauss Summation
The Gauss batch begins by generating 363 conjectures in 106 seconds — the larger conjecture creation time reflecting the richer definition set, which now includes intervals, sequences, limit constructs, and the fold operator. Before the prover starts, the anchor connection mechanism automatically attaches the Gauss anchor to the Peano anchor, injecting the implication AnchorGauss $\Rightarrow$ AnchorPeano into the theorem list. This ensures that all 64 theorems proven in Batch 1 are available as ground knowledge for the Gauss prover.
The CE filter runs for 192.5 seconds — the dominant cost in this batch — passing 29 conjectures to the prover. At the start of Phase 2, the prover broadcasts the 64 inherited Peano theorems across all 70 active memory blocks, making prior knowledge immediately available to every Logic Block. The prover then runs 30 iterations (2 warm-up, 28 main) and finishes in 47.1 seconds, proving 10 theorems including the flagship result: Gauss’s summation formula. The total runtime of the native executable for this batch is 239.7 seconds.
@@ 4 Case Study: Elementary Number Theory
Table 1: End-to-end runtime by phase and batch on commodity hardware (Dell G16 7630, 32 logical cores).
Phase Peano (s) Share (%) Gauss (s) Share (%)
Conjecturer 21.3 5.2 106.0 26.0
CE filter 14.6 3.6 192.5 47.3
Prover 24.6 6.0 47.1 11.6
Total (executable) 39.4 9.7 239.7 58.8
Overall wall-clock: 407 seconds
Observations
The runtime distribution shifts considerably between batches. In the Peano batch the three phases are roughly balanced; in the Gauss batch the CE filter becomes the dominant cost at 80% of the executable runtime, reflecting the increased complexity of the fact base required to screen a richer conjecture space. The prover, by contrast, remains tractable at under 50 seconds even with 64 inherited theorems broadcast into the grid. Conjecture creation accounts for the bulk of the Python-side runtime and remains the primary target for the planned migration to C++. These figures should be read as a snapshot of a initial prototype prototype; substantial headroom remains across all phases.
@@ 5 Related Work and Contributions
Systems such as Lean, Rocq/Coq, and Isabelle/HOL provide expressive foundations for formalizing mathematics and verifying proofs in a small, trusted kernel [6]. Although their traditional workflow is interactive—users decompose goals, invoke tactics, and supply key intermediate lemmas—modern ecosystems expose substantial automation. Isabelle’s Sledgehammer tool, for example, exports the current goal together with a heuristic relevance slice of the local theory context to a portfolio of external automated theorem provers (ATPs) and SMT solvers (e.g., E, Vampire, Z3, cvc5); when one succeeds, Isabelle attempts proof reconstruction automatically [3]. Lean and Coq likewise support tactic automation and scripted, batch replay of large libraries. In short, contemporary ITPs can discharge many individual goals automatically once the goals and supporting libraries are in place. Coq’s ecosystem has recently incorporated “hammer” plugins (e.g. CoqHammer) that automatically attempt to prove goals by translating them to first-order logic and calling external provers [4] , often discharging nontrivial goals without manual guidance. Automated first-order provers like E have achieved remarkable success in solving individual conjectures, consistently ranking among the top systems in the annual CASC competitions[10]. Another premier first-order prover is Vampire, which has dominated many competition divisions – since 1999 it has won over 50 category titles in the “world cup” of theorem provers [7] – thanks to its highly optimized saturation engine. Satisfiability Modulo Theories (SMTs) solvers (e.g. Z3) excel at deciding logical formulas in rich theories automatically [5] , and they have become integral to many verification workflows (often being invoked behind the scenes by proof assistant tactics to discharge specific subgoals). The newest SMT solvers, such as cvc5, extend the capabilities of their predecessors (e.g. CVC4 and Z3) with support for a diverse range of theories and advanced features like higher-order reasoning and syntax-guided synthesis [2].
@@ 6 Future Work, Roadmap and Potential Impact
With the Gauss summation formula proven and shipped, GL has completed its first two-batch ENT run. Peano delivered algebraic manipulation; Gauss delivered logical transformation. That closes one phase. The next milestone is the Fundamental Theorem of Arithmetic (FTA).
The road to FTA is not a continuation of the same work at larger scale. It requires three infrastructure extensions and one final major algorithmic extension, arriving in a specific order dictated by their dependencies.
6.1 Pruner
GL now carries a growing mass of proven theorems. That is good for expressive power but bad for focus. The pruner’s task is to select a minimal core subset from that mass such that all remaining theorems follow directly from it. This is not merely a runtime optimization. It is a structural compression step: it removes noise from theorem selection, improves proof trace readability, and ensures that the HTML proof graph reflects the real backbone of a result rather than drowning it in nearby consequences. As the theorem base grows, this compression step becomes essential for maintaining clarity and performance.
6.2 External Proof Checker
Internal proof generation is not sufficient once GL begins carrying heavier mathematical claims. An external checker creates a hard boundary between proving and verifying. GL generates proof objects; the checker validates them independently, without access to the internal machinery of the prover. This improves trust, makes debugging cleaner, and gives the project a stronger research posture. When the prover is modified, the checker remains the stable witness. When something fails, the boundary makes it immediately clear whether the issue lies in generation or in verification.
6.3 Case Differentiation
This is the last missing tool in GL’s mathematical stack, and arguably the most consequential. Peano gave GL algebraic manipulation. Gauss gave GL logical transformation. Case differentiation is the third pillar.
A significant portion of non-trivial mathematics is not blocked by algebra or by rewrite logic — it is blocked by the need to split on a condition and drive separate proof branches under separate assumptions. GL currently follows a single linear path through a proof. Once case differentiation is in place, the system can handle proofs that depend on sign splits, divisibility splits, structural alternatives, and contradiction-driven branch closures. That is the point at which the prover stops being a corridor and becomes a machine capable of operating in branching terrain.
6.4 Incubator
The Incubator becomes possible once case differentiation is available, because it depends on that tool to generate its outputs reliably. The incubator autogenerates ground-level fact tables — the machine-generated equivalent of multiplication tables — and feeds them into the CE filter as fuel. The CE filter grows stronger in proportion to the richness of its fact stock. The incubator is therefore a throughput engine: it manufactures the raw material that allows the CE filter to operate at higher quality and higher speed, without requiring manual curation of facts as the target domain grows.
6.5 The FTA Campaign
Once the above stack is complete, the FTA campaign becomes a direct engineering effort. A route of auxiliary facts and intermediate theorems leading from the Peano axioms toward FTA will be identified and pressure-tested in sequence. If the conjecturer fails to surface a required fact, the conjecturer is extended. If the prover cannot establish it, the prover is extended. That loop repeats until the route is carried end to end.
6.6 Path to Scalability
In parallel with the mathematical extensions above, the implementation will be progressively hardened. The Python conjecturer — currently the dominant runtime cost — will be migrated to C++, recovering the multithreading headroom that Python’s GIL forecloses. The architecture will then be adapted for massively parallel cloud execution, distributing the workload across large numbers of instances on platforms such as AWS or GCP. The long-term target substrate remains a cloud of specialized ASIC chips, where the full parallelism of the Logic Block grid can be realized without constraint.